OpenStudy (kaloemalia):

PLEASE HELP 2 1/2 times 8/9

jimthompson5910 (jim_thompson5910):

are you able to convert `2 1/2` to an improper fraction?

OpenStudy (kaloemalia):

yes

OpenStudy (kaloemalia):

I did that

jimthompson5910 (jim_thompson5910):

what is `2 1/2` equal to

OpenStudy (kaloemalia):

5/2

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so you really have 5/2 times 8/9

OpenStudy (kaloemalia):

yes.

jimthompson5910 (jim_thompson5910):

multiply straight across (the numerators pair up and multiply; so do the denominators)

jimthompson5910 (jim_thompson5910):

after multiplying, reduce as much as possible

OpenStudy (kaloemalia):

40/18? then reduce right?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (kaloemalia):

20/9

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

optionally you can convert that result to a mixed number

OpenStudy (kaloemalia):

but when i tried to turn back into a mixed number but it is a never ending 2.2222222222

jimthompson5910 (jim_thompson5910):

break that decimal into 2 parts the whole part which is 2 (left of the decimal) the fractional part which is 0.2222 repeating (right of the decimal)

jimthompson5910 (jim_thompson5910):

the decimal part 0.2222 gets multiplied with the denominator of 9 to get roughly 0.22222*9 = 1.99998 which is effectively 2 (there's a bit of rounding error) so 2.222222... = 2 2/9

jimthompson5910 (jim_thompson5910):

if you had something like 3.2222222..., then 3.2222222... = 3 2/9 or if you had 7.2222222..., then 7.2222222... = 7 2/9 and so on

jimthompson5910 (jim_thompson5910):

or you can think of it like this \[\Large \frac{20}{9} = \color{red}{2} \text{ remainder } \color{blue}{2}\] \[\Large \frac{20}{9} = \color{red}{2} \frac{\color{blue}{2}}{9}\]

OpenStudy (kaloemalia):

so 2 2/9 is my final answer?

jimthompson5910 (jim_thompson5910):

yes if they want the answer as a mixed number

OpenStudy (kaloemalia):

okay! Thank you so much!

jimthompson5910 (jim_thompson5910):

glad to be of help

OpenStudy (mathstudent55):